# Adversarial Robustness in Parameter-Space Classifiers

### v1

@@ B.1.1 MNIST
Figure 6 illustrates the mean amplification difference measured through model activation layers over MNIST data, similarly to Figure 5, however for the signal-domain classifier trained according to section A.3.2. This model uses a 3-layer convolutional encoder, followed by a classifier with a single activation. Clean classifier accuracy is $98.8\%$, robust accuracy is $11.28\%$. Differently from similar results displayed for parameter-space classifiers in Figure 5, upon transition from encoder to classifier layers (3rd layer, marked by the black vertical line), the adversarial amplification is not attenuated, but rather drastically inflated.
@@
Figure 7 showcases adversarial amplification (similarly to Figure 5) across different attack bound for the full-PGD attack. Adversarial obfuscation drawn from Figure 5 arises here as well.
@@ B.1.2 Fashion-MNIST
Figure 8 illustrates adversarial amplification across pipeline layers, similarly to Figure 5. Trends of correlation between amplification and adversarial success, as well as evident obfuscation of adversarial affect between INR and classifier segments, are consistent with findings from Section 4.3.
@@
Similarly to Figure 6, we test amplification patterns of a signal-domain classifier (Section A.3.2) for the Fashion-MNIST dataset as well. We again observe consistency across both datasets - the signal domain classifiers proves to be much more sensitive to adversarial examples propagated through the encoder-classifier pipeline.
@@
Figure 10 exhibits the affect of adversarial attacks in signal-domain over the underlying modulation vectors (for parameter-space classifiers) and latent vectors (for signal-space classifiers) received as classifier input. Similarly to Figure 1, for both parameter-space and signal-space model, a well-formed latent space is observed before the incorporation of adversarial perturbation. After attacking with PGD, the corresponding modulation vectors fed into the parameter-space classifier remain almost unchanged, whilst the latent space observed by the signal-domain classifier is heavily deformed.
@@ B.2 Robustness For 3D Data
In Table 2 we report numerical results for the data from Figure 4 The first two columns represent the same attack constraints in absolute bit numbers ($\varepsilon$) and bit numbers relative to the entire volume (matching horizontal axis in figure 4).

### v2

@@ B.1.1 MNIST
fig. 5 illustrates the mean amplification difference measured through model activation layers over MNIST data, similarly to fig. 3, however for the signal-domain classifier trained according to section A.4.2. This model uses a 3-layer convolutional encoder, followed by a classifier with a single activation. Clean classifier accuracy is $98.8\%$, robust accuracy is $11.28\%$. Differently from similar results displayed for parameter-space classifiers in fig. 3, upon transition from encoder to classifier layers (3rd layer, marked by the black vertical line), the adversarial amplification is not attenuated, but rather drastically inflated.
@@
fig. 6 showcases adversarial amplification (similarly to fig. 3) across different attack bound for the full-PGD attack. Adversarial obfuscation drawn from fig. 3 arises here as well.
@@ B.1.2 Fashion-MNIST
fig. 7 illustrates adversarial amplification across pipeline layers, similarly to fig. 3. Trends of correlation between amplification and adversarial success, as well as evident obfuscation of adversarial affect between INR and classifier segments, are consistent with findings from section 4.3.
@@
Similarly to fig. 5, we test amplification patterns of a signal-domain classifier (section A.4.2) for the Fashion-MNIST dataset as well. We again observe consistency across both datasets - the signal domain classifiers proves to be much more sensitive to adversarial examples propagated through the encoder-classifier pipeline.
@@
fig. 9 exhibits the affect of adversarial attacks in signal-domain over the underlying modulation vectors (for parameter-space classifiers) and latent vectors (for signal-space classifiers) received as classifier input. Similarly to fig. 4, for both parameter-space and signal-space model, a well-formed latent space is observed before the incorporation of adversarial perturbation. After attacking with PGD, the corresponding modulation vectors fed into the parameter-space classifier remain almost unchanged, whilst the latent space observed by the signal-domain classifier is heavily deformed.
@@ B.2 Robustness For 3D Data
In table 2 we report numerical results for the data from fig. 2(b) The first two columns represent the same attack constraints in absolute bit numbers ($\varepsilon$) and bit numbers relative to the entire volume (matching horizontal axis in figure 2(b)).
@@
Appendix C Computational Resource Analysis
In sections 4.1,4.2 we provide evidence for the adversarial robustness of parameter-space classifiers based on classification accuracies under the presence of adversarial perturbations. In this section, we further support these claims from a different, practical perspective. We show that orthogonally to the efficacy of adversarial attacks over parameter-space classifiers, attacking these classifiers also poses significant added computational difficulties for the potential attacker. These difficulties stem from the optimization process inherent within the classification pipeline’s inference process. To illustrate this, we measure the computation times required for a single modulation optimization step during a single PGD iteration on an NVIDIA RTX-2080 GPU. Our results show that for the MNIST dataset with 512-dimensional modulation vectors, the compute time per step is approximately 0.015 seconds, while for the ModelNet10 dataset with 2048-dimensional modulation vectors, it is 0.02 seconds. Even for a small number of PGD steps, increasing the number of modulation optimization steps – an aspect controlled by the model, not the attacker – increases attack optimization times within several orders of magnitude. Importantly, this increase does not significantly impact the computation time required for ”clean” (unperturbed) inference. For instance, with 100 modulation optimization steps set by the model to solve eq. 1, ”clean” inference would require 1.5 seconds per sample, while perturbation optimization of 100 PGD iterations would require the attacker to invest 150 seconds per-sample.